OpenStudy (bananas1234):

Simplify: -27a^2b^2/9ab

OpenStudy (mrnood):

just cancel the 9 with the -27 a with a^2 b with b^2

OpenStudy (mrnood):

This is the same as \[\frac{ -27 }{ 9 } \times \frac{ a ^{2} }{ a } \times \frac{ b ^{2} }{ b }\]